OpenStudy (anonymous):

A motorboat travels 522 in 6 hours going upstream and 882 in 7 hours going downstream. What is the rate of the boat in still water and what is the rate of the current?

OpenStudy (lgbasallote):

you're missing some units of measurement in 522 and 882

OpenStudy (anonymous):

miles

OpenStudy (anonymous):

sorry

OpenStudy (lgbasallote):

hmm ok...you know that speed = distance/time right?

OpenStudy (anonymous):

i did not

OpenStudy (lgbasallote):

well now you do

OpenStudy (lgbasallote):

one more thing... downstream speed = speed of boat + speed of current upstream speed = speed of boat - speed of current

OpenStudy (lgbasallote):

got that?

OpenStudy (anonymous):

uhh maybe ill try plug in the numbers just check my work

OpenStudy (lgbasallote):

okay then

OpenStudy (anonymous):

speed of current?

OpenStudy (lgbasallote):

what do you mean?

OpenStudy (anonymous):

i have know idea how to get that

OpenStudy (lgbasallote):

okay then..shall i step in?

OpenStudy (anonymous):

i think that make me grasp this problem

OpenStudy (anonymous):

would*

OpenStudy (lgbasallote):

okay then...first let's use letters to make these shorter let speed of boat = b speed of current = c therefore downstream speed = b + c upstream speed = b - c okay?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

upstream speed 87

OpenStudy (lgbasallote):

btw...can you recheck your question? is it really like that?

OpenStudy (lgbasallote):

just asking

OpenStudy (anonymous):

yes it wants me to find the rate boats in still water and rate of the current

OpenStudy (lgbasallote):

ok good...so you're right upstream speed is 87 what is downstream speed then?

OpenStudy (anonymous):

126

OpenStudy (lgbasallote):

good..let's plug those in 126 = b + c 87 = b - c are you seeing the next step now? :)

OpenStudy (anonymous):

sadly no

OpenStudy (lgbasallote):

you use systems of equations ;) see how if you add the equations c will cancel out?

OpenStudy (anonymous):

so 126c=b

OpenStudy (lgbasallote):

hmm that's not what i meant hehe

OpenStudy (anonymous):

im sorry

OpenStudy (lgbasallote):

but we can use substitution..but it would be b = 126 - c got it?

OpenStudy (anonymous):

ight now what

OpenStudy (lgbasallote):

now substitute that into the b of 87 = b - c it will ebcome 87 = 126 - c - c got it?

OpenStudy (anonymous):

c=39

OpenStudy (lgbasallote):

not exactly... 87 - 126 = -c - c -39 = -2c

OpenStudy (lgbasallote):

you have to find c

OpenStudy (anonymous):

c=19.5

OpenStudy (lgbasallote):

yup better

OpenStudy (anonymous):

so?

OpenStudy (lgbasallote):

now that is the speed of current with this you can also find the speed of the boat

OpenStudy (anonymous):

how bro? if havent caught im kinda bad at math

OpenStudy (lgbasallote):

nahh you caught up fast remember how 126 = b + c

OpenStudy (anonymous):

so the rate of the current is 19.5

OpenStudy (lgbasallote):

yup

OpenStudy (anonymous):

b=106.5

OpenStudy (anonymous):

?

OpenStudy (lgbasallote):

yup

OpenStudy (lgbasallote):

you can check by seeing if 87 = 106.5 - 19.5

OpenStudy (anonymous):

so the rate of th boat in still water is b?